OpenStudy (anonymous):

[HOW THIS EQUAL?] -2(e^t)(e^-t)+e^(-2t)+4=e^(2t)+2+e^(-2t)

OpenStudy (anonymous):

\[-2e^te^{-t}+e^{-2t}+4=e ^{2t}+2+e ^{-2t}\]

OpenStudy (anonymous):

how do i simply simplify this....?

OpenStudy (anonymous):

\[-2e^te^{-t}+e^{-2t}+4=e ^{2t}+2+e ^{-2t}\]

OpenStudy (anonymous):

wait sorry i put wrong..

OpenStudy (anonymous):

-2e^te^{-t}+e^{-2t}+4= 2+e ^{-2t}

OpenStudy (anonymous):

\[-2e^te^{-t}+e^{-2t}+4= 2+e ^{-2t}\]

OpenStudy (anonymous):

wait lemme close and re open it